OpenStudy (anonymous):

HELP PLEASE!! The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, −2). Answers: y = 2x + 4 y = negative 1 over 2x − 1 over 2 y = − 1 over 2x − 7 over 2 y = 2x − 8

OpenStudy (anonymous):

step one find the slope (with your eyeballs) of \( y = 2x + 4\)

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

well close, it is the number \(2\) without the variable

OpenStudy (anonymous):

oh ok well 2

OpenStudy (anonymous):

step 2 know that a parallel line has the same slope, so the slope of the line you are looking for is also \(2\)

OpenStudy (anonymous):

so that leaves us with only answers a & b?

OpenStudy (anonymous):

step3 write the points slope formula \(y-y_1=m(x-x_1)\) with \[m=2,x_1=3,y_1=-2\]

OpenStudy (anonymous):

actually leaves A and D right?

OpenStudy (anonymous):

yes my mistake hold on im solving it now!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

y-2=2(x-3), so y-2= 2x-6

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

oh now i add the 2 to the other side to cancel it out

OpenStudy (anonymous):

\(y_1=-2\) so \[y-y_1=y+2\] on the left

OpenStudy (anonymous):

so y=2x-8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw hope it is more or less clear